### v1

@@ 3.5 Probabilistic bounds
Let $\mathcal{A}:L^{2}(D)\to L^{2}(D)$ be a non-negative self-adjoint trace-class operator, $2\leq k\leq\operatorname{rank}(\mathcal{A})$ be a target rank, and $p\geq 4$ be an oversampling parameter. Let $\Omega$ be a quasimatrix with $k+p$ columns i.i.d. from $\mathcal{GP}(0,K)$, with a kernel $K$ such that the matrix $\widetilde{\bm{K}}_{11}$ defined in Eq. 3 is invertible. Then, the Nyström approximation $\widehat{\mathcal{A}}:L^{2}(D)\to L^{2}(D)$ defined in Eq. 20 satisfies
$\mathbb{E}[\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{HS}}]$ $\leq\left(1+2\delta_{k}^{(\mathrm{HS})}+2\sqrt{c_{1}}\beta_{k}^{(\mathrm{HS})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+2\sqrt{c_{2}}\beta_{k}^{(\mathrm{Tr})}\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}},$ (23a)
$\mathbb{E}[\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{op}}]$ $\leq\left(1+\frac{3k}{p-1}\beta_{k}^{(\mathrm{op})}+3\delta_{k}^{(\mathrm{op})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{op}}+\frac{3e^{2}(k+p)}{p^{2}-1}\beta_{k}^{(\mathrm{Tr})}\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}},$ (23b)
$\mathbb{E}[\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{Tr}}]$ $\leq\left(1+\frac{k}{p-1}\beta_{k}^{(\mathrm{Tr})}+\delta_{k}^{(\mathrm{Tr})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}},$ (23c)
@@
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{HS}}$ $\leq\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+4\left(\delta_{k}^{(\mathrm{HS})}+t^{2}(d_{1}+d_{3})\beta_{k}^{(\mathrm{HS})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+4t^{2}d_{3}\beta_{k}^{(\mathrm{Tr})}\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}}$ (24a)
$\quad+2t^{2}u^{2}d_{2}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{\mathrm{op}},$
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{op}}$ $\leq\left(1+4\delta_{k}^{(\mathrm{op})}+4(d_{1}+d_{2}u^{2})t^{2}\beta_{k}^{(\mathrm{op})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{op}}+4d_{2}t^{2}\beta_{k}^{(\mathrm{Tr})}\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}},$ (24b)
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{Tr}}$ $\leq\left(1+2\delta_{k}^{(\mathrm{Tr})}+d_{1}t^{2}\beta_{k}^{(\mathrm{Tr})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}}+2d_{2}t^{2}u^{2}\beta_{k}^{(\mathrm{op})}\|\bm{\Sigma}_{2}\|_{\mathrm{op}},$ (24c)
with probability $\geq 1-3t^{-p}-e^{-u^{2}/2}$. Here, $d_{1}=\mathcal{O}(k/p)$, $d_{2}=\mathcal{O}(k/p)$, and $d_{3}=\mathcal{O}(k^{3/2}/p)$ are the constants defined in Eq. 12.
@@
$X_{s,n}=\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Omega}_{2})_{\llbracket 1,n\rrbracket\times\llbracket 1,k+p\rrbracket}\bm{\Omega}_{1}^{\dagger}\|_{(2s)},\quad X_{s}=\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(2s)},$ (25)
@@
For $s\in[1,\infty]$, let $X_{s}$ be the random variable defined in Eq. 25. Then, if $p\geq 4$, we have $\mathbb{E}_{\Omega}^{2}[X_{s}]<\infty$.
Proof 3.6.
@@
For $s\in[1,\infty]$, let $X_{s}$, $X_{s,n}$ be the random variables defined in Eq. 25 for $n\geq 1$. Then, if $p\geq 4$, we have $\lim_{n\rightarrow\infty}\mathbb{E}^{2}[X_{s}-X_{s,n}]=0$.
Proof 3.8.
For $n\geq 1$, we define the quasimatrix $\bm{\Omega}_{2}^{(n)}$ whose first $n$ rows are equal to the first $n$ rows of $\bm{\Omega}_{2}$ and the remaining rows are zero. Then,
@@
[…]
@@
Following [8, Sec. 3.3], we know that the entries of $\bm{\omega}_{1}$ satisfy $(\bm{\omega}_{1})_{i}\sim\mathcal{N}(0,(\widetilde{\bm{K}}_{22})_{ii})$ for $i\in\mathbb{N}^{*}$. Let $Y=\sum_{i=n+1}^{\infty}(\bm{\omega}_{1})_{i}^{2}=\|\bm{\omega}_{1}-\bm{\omega}_{1}^{(n)}\|_{2}^{2}$. Combining the non-negativity of the summands and the Fubini–Tonelli theorem, we can interchange the summation and expectation to obtain
@@
where the last inequality follows from the Cauchy–Schwarz inequality. Hence,
@@
Combining Lemmas 3.5 and 3.7, we obtain that $\lim_{n\rightarrow\infty}\mathbb{E}^{2}[X_{s,n}]=\mathbb{E}^{2}[X_{s}]$. Hence, if we have a family of bounds $\mathbb{E}^{2}[X_{s,n}]\leq y_{s,n}$ with $y_{s,n}\to y_{s}$ as $n\to\infty$, then $\mathbb{E}^{2}[X_{s}]\leq y_{s}$. Furthermore, combining the continuous mapping theorem and the fact that $L^{2}$ convergence implies convergence in distribution, for any positive sequence $z_{s,n}\to z_{s}<\infty$, we have $\lim_{n\rightarrow\infty}\mathbb{P}(X_{s,n}^{2}>z_{s,n})=\mathbb{P}(X_{s}^{2}>z_{s})$. We can now proceed with the proof of Theorem 3.4, which uses results from Section 2 to derive expressions for $y_{s,n}$ and $z_{s,n}$ and show that, for $s\in\{2,\infty,1\}$, they converge to the right-hand sides of Theorems 3.4 and 3.4.
Proof 3.9 (Proof of Theorem 3.4).
Let $\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+y_{2}$, $\|\bm{\Sigma}_{2}\|_{\mathrm{op}}+y_{\infty}$, and $\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}}+y_{1}$ be the respective right-hand sides in Theorem 3.4, and $\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+z_{2}$, $\|\bm{\Sigma}_{2}\|_{\mathrm{op}}+z_{\infty}$, and $\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}}+z_{1}$ be the right-hand sides in Theorem 3.4. For $s\in\{2,\infty,1\}$, we aim to prove that $\mathbb{E}[X_{s}^{2}]\leq y_{s}$ and $\mathbb{P}(X_{s}^{2}\geq z_{s})\leq 3t^{-p}+e^{-u^{2}/2}$. Following Theorem 2.1, we have
$\mathbb{E}[X_{2,n}^{2}]$ $\leq 2\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}\widetilde{\bm{K}}_{11}^{-1}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}^{*}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\textup{F}}$
$\quad+2\sqrt{c_{1}}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\textup{F}}$
$\quad+2\sqrt{c_{2}}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{*}:=y_{2,n},$
@@
$X_{2,n}^{2}$ $\geq 4\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}\widetilde{\bm{K}}_{11}^{-1}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}^{*}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\textup{F}}$
$\quad+4t^{2}(d_{1}+d_{3})\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\textup{F}}$
$\quad+4t^{2}d_{3}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{*}$
$\quad+2t^{2}u^{2}d_{2}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{2}:=z_{2,n}.$
Following Theorems 2.4 and 2.6 we know that
@@
where $y_{\infty,n}$, $y_{1,n}$, $z_{\infty,n}$, and $z_{1,n}$ can be defined analogously to $y_{2,n}$ and $z_{2,n}$ using Theorems 2.4 and 2.6. Moreover, Lemma 3.7 implies that $\lim_{n\rightarrow\infty}\mathbb{E}[X_{s,n}^{2}]=\mathbb{E}[X_{s}^{2}]$ and convergence in distribution, which implies $\lim_{n\rightarrow\infty}\mathbb{P}(X_{s,n}>z_{s,n})=\mathbb{P}(X_{s}>\lim_{n\rightarrow\infty}z_{s,n})$. Hence, it is sufficient to show that $\lim_{n\rightarrow\infty}y_{s,n}=y_{s}$ and $\lim_{n\rightarrow\infty}z_{s,n}=z_{s}$ for $s\in\{2,\infty,1\}$. For this purpose, it is sufficient to show that
@@
$\lim_{n\rightarrow\infty}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}\widetilde{\bm{K}}_{11}^{-1}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}^{*}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\xi}=$ (26)
$\hskip 273.18271pt\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2}\|_{\xi},$
@@
$\hskip 260.17464pt(\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}.$
Since $\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}\bm{\Sigma}_{2}^{1/2}$ and $\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2}$ are non-negative trace-class operators, applying (22a) yields (26), as desired.

### v2

@@ 3.5 Probabilistic bounds
Let $\mathcal{A}:L^{2}(D)\to L^{2}(D)$ be a non-negative self-adjoint trace-class operator, $2\leq k\leq\operatorname{rank}(\mathcal{A})$ be a target rank, and $p\geq 4$ be an oversampling parameter. Let $\Omega$ be a quasimatrix with $k+p$ columns i.i.d. from $\mathcal{GP}(0,K)$, with a kernel $K$ such that the matrix $\widetilde{\bm{K}}_{11}$ defined in (3) is invertible. Then, the Nyström approximation $\widehat{\mathcal{A}}:L^{2}(D)\to L^{2}(D)$ defined in Eq. 23 satisfies
(26a) $\mathbb{E}[\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{op}}]$ $\leq\left(1+\frac{3k}{p-1}\beta_{k}^{(\mathrm{op})}+3\delta_{k}^{(\mathrm{op})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{op}}+\frac{3e^{2}(k+p)}{p^{2}-1}\beta_{k}^{(\mathrm{Tr})}\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}},$
(26b) $\mathbb{E}[\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{Tr}}]$ $\leq\left(1+\frac{k}{p-1}\beta_{k}^{(\mathrm{Tr})}+\delta_{k}^{(\mathrm{Tr})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}},$
(26c) ${\color[rgb]{0,0,0}\mathbb{E}[\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{HS}}]}$ ${\color[rgb]{0,0,0}\leq\left(1+2\delta_{k}^{(\mathrm{HS})}+2\sqrt{c_{1}}\beta_{k}^{(\mathrm{HS})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+2\sqrt{c_{2}}\beta_{k}^{(\mathrm{Tr})}\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}},}$
@@
(27a) $\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{op}}$ $\leq\left(1+4\delta_{k}^{(\mathrm{op})}+4(d_{1}+d_{2}u^{2})t^{2}\beta_{k}^{(\mathrm{op})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{op}}+4d_{2}t^{2}\beta_{k}^{(\mathrm{Tr})}\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}},$
(27b) $\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{Tr}}$ $\leq\left(1+2\delta_{k}^{(\mathrm{Tr})}+d_{1}t^{2}\beta_{k}^{(\mathrm{Tr})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}}+2d_{2}t^{2}u^{2}\beta_{k}^{(\mathrm{op})}\|\bm{\Sigma}_{2}\|_{\mathrm{op}},$
(27c) ${\color[rgb]{0,0,0}\|\mathcal{A}-\widehat{\mathcal{A}}\|_{\mathrm{HS}}}$ ${\color[rgb]{0,0,0}\leq\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+4\left(\delta_{k}^{(\mathrm{HS})}+t^{2}(d_{1}+d_{3})\beta_{k}^{(\mathrm{HS})}\right)\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+4t^{2}d_{3}\beta_{k}^{(\mathrm{Tr})}\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}}}$
${\color[rgb]{0,0,0}\quad+2t^{2}u^{2}d_{2}\beta_{k}^{(2)}\|\bm{\Sigma}_{2}\|_{\mathrm{op}},}$
with probability $\geq 1-3t^{-p}-e^{-u^{2}/2}$. Here, $d_{1}=\mathcal{O}(k/p)$, $d_{2}=\mathcal{O}(k/p)$, and $d_{3}=\mathcal{O}(k^{3/2}/p)$ are the constants defined in Eq. 7 and in Theorem 2.10.
@@
(28) $X_{s,n}=\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Omega}_{2})_{\llbracket 1,n\rrbracket\times\llbracket 1,k+p\rrbracket}\bm{\Omega}_{1}^{\dagger}\|_{(2s)},\quad X_{s}=\|\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\|_{(2s)},$
@@
For $s\in[1,\infty]$, let $X_{s}$ be the random variable defined in (28). Then, if $p\geq 4$, we have $\mathbb{E}_{\Omega}^{2}[X_{s}]<\infty$.
Proof 3.7.
@@
For $s\in[1,\infty]$, let $X_{s}$, $X_{s,n}$ be the random variables defined in (28) for $n\geq 1$. Then, if $p\geq 4$, we have $\lim_{n\rightarrow\infty}\mathbb{E}^{2}[X_{s}-X_{s,n}]=0$.
Proof 3.9.
For $n\geq 1$, we define the matrix $\bm{\Omega}_{2}^{(n)}$, with an infinite number of rows, whose first $n$ rows are equal to the first $n$ rows of $\bm{\Omega}_{2}$ and the remaining rows are zero. Then,
@@
By Hölder’s inequality it suffices to show that $\lim_{n\rightarrow\infty}\mathbb{E}[\|\bm{\Omega}_{2}-\bm{\Omega}_{2}^{(n)}\|_{\mathrm{HS}}^{4}]=0$, since $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}]<\infty$ by Lemma A.3. Let $\bm{\omega}_{i}$ and $\bm{\omega}_{i}^{(n)}$ denote the $i^{\text{th}}$ columns of $\bm{\Omega}_{2}$ and $\bm{\Omega}_{2}^{(n)}$, respectively. Using the monotonicity and triangle inequality of $L^{p}$-norms, we have
@@
Following [8, Sec. 3.3], we know that the entries of $\bm{\omega}_{1}$ satisfy $(\bm{\omega}_{1})_{i}\sim\mathcal{N}(0,(\widetilde{\bm{K}}_{22})_{ii})$ for $i\in\mathbb{N}^{*}$. Let $Y=\sum_{i=n+1}^{\infty}(\bm{\omega}_{1})_{i}^{2}=\|\bm{\omega}_{1}-\bm{\omega}_{1}^{(n)}\|_{2}^{2}$. Combining the non-negativity of the summands and the Fubini-Tonelli theorem, we can interchange the summation and expectation to obtain
@@
where the last inequality follows from the Cauchy-Schwarz inequality. Hence,
@@
Combining Lemmas 3.6 and 3.8, we obtain that $\lim_{n\rightarrow\infty}\mathbb{E}^{2}[X_{s,n}]=\mathbb{E}^{2}[X_{s}]$. Hence, if we have a family of bounds $\mathbb{E}^{2}[X_{s,n}]\leq y_{s,n}$ with $y_{s,n}\to y_{s}$ as $n\to\infty$, then $\mathbb{E}^{2}[X_{s}]\leq y_{s}$. Furthermore, combining the continuous mapping theorem and the fact that $L_{2}$-convergence implies convergence in distribution, for any positive sequence $z_{s,n}\to z_{s}<\infty$, we have $\lim_{n\rightarrow\infty}\mathbb{P}(X_{s,n}^{2}>z_{s,n})=\mathbb{P}(X_{s}^{2}>z_{s})$. We can now proceed with the proof of Theorem 3.5, which uses results from Section 2 to derive expressions for $y_{s,n}$ and $z_{s,n}$ and show that, for $s\in\{2,\infty,1\}$, they converge to the right-hand sides of Eqs. 26 and 27.
Proof 3.10 (Proof of Theorem 3.5).
Let $\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+y_{2}$, $\|\bm{\Sigma}_{2}\|_{\mathrm{op}}+y_{\infty}$, and $\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}}+y_{1}$ be the respective right-hand sides in Eq. 26, and $\|\bm{\Sigma}_{2}\|_{\mathrm{HS}}+z_{2}$, $\|\bm{\Sigma}_{2}\|_{\mathrm{op}}+z_{\infty}$, and $\|\bm{\Sigma}_{2}\|_{\mathrm{Tr}}+z_{1}$ be the right-hand sides in Eq. 27. For $s\in\{2,\infty,1\}$, we aim to prove that $\mathbb{E}[X_{s}^{2}]\leq y_{s}$ and $\mathbb{P}(X_{s}^{2}\geq z_{s})\leq 3t^{-p}+e^{-u^{2}/2}$. Following Theorem 2.8, we have
$\mathbb{E}[X_{2,n}^{2}]$ $\leq 2\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}{\color[rgb]{0,0,0}\widetilde{\bm{K}}_{11}^{-2}}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}^{*}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\textup{F}}$
$\quad+2\sqrt{c_{1}}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\textup{F}}{\color[rgb]{0,0,0}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}$
$\quad+2\sqrt{c_{2}}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{*}{\color[rgb]{0,0,0}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}:=y_{2,n},$
@@
$X_{2,n}^{2}$ $\geq 4\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}{\color[rgb]{0,0,0}\widetilde{\bm{K}}_{11}^{-2}}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}^{*}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\textup{F}}$
$\quad+4t^{2}(d_{1}+d_{3})\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\textup{F}}{\color[rgb]{0,0,0}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}$
$\quad+4t^{2}d_{3}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{*}{\color[rgb]{0,0,0}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}$
$\quad+2t^{2}u^{2}d_{2}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{22.1})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{2}{\color[rgb]{0,0,0}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}:=z_{2,n}.$
Following Theorems 2.1 and 2.5 we know that
@@
where $y_{\infty,n}$, $y_{1,n}$, $z_{\infty,n}$, and $z_{1,n}$ can be defined analogously to $y_{2,n}$ and $z_{2,n}$ using Eqs. 6 and 8. Moreover, Lemma 3.8 implies that $\lim_{n\rightarrow\infty}\mathbb{E}[X_{s,n}^{2}]=\mathbb{E}[X_{s}^{2}]$ and convergence in distribution, which implies $\lim_{n\rightarrow\infty}\mathbb{P}(X_{s,n}>z_{s,n})=\mathbb{P}(X_{s}>\lim_{n\rightarrow\infty}z_{s,n})$. Hence, it is sufficient to show that $\lim_{n\rightarrow\infty}y_{s,n}=y_{s}$ and $\lim_{n\rightarrow\infty}z_{s,n}=z_{s}$ for $s\in\{2,\infty,1\}$. For this purpose, it is sufficient to show that
@@
(29) $\lim_{n\rightarrow\infty}\|(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}{\color[rgb]{0,0,0}\widetilde{\bm{K}}_{11}^{-2}}(\widetilde{\bm{K}}_{21})_{\llbracket 1,n\rrbracket\times\llbracket 1,k\rrbracket}^{*}(\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\xi}=$
$\hskip 273.18271pt\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}{\color[rgb]{0,0,0}\widetilde{\bm{K}}_{11}^{-2}}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2}\|_{\xi},$
@@
$\hskip 260.17464pt(\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}{\color[rgb]{0,0,0}\widetilde{\bm{K}}_{11}^{-2}}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2})_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}.$
Since $\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}\bm{\Sigma}_{2}^{1/2}$ and $\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}{\color[rgb]{0,0,0}\widetilde{\bm{K}}_{11}^{-2}}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2}$ are non-negative trace-class operators, applying (25a) yields (29), as desired.